OpenStudy (anonymous):

use logarithmic differentiation to find the derivative of y with respect to x:

OpenStudy (anonymous):

\[y=\sqrt[3]{\frac{ x(x+1)(x-2) }{ (x ^{2}+1)(2x+3)}}\]

OpenStudy (anonymous):

ln[y] = (1/3) ln [ (x(x+1)(x-2)) / ((x^2+1)(2x+3)) ] take the derivative now : (dy/dx) / y = (1/3) * [ (x(x+1)(x-2)) / ((x^2+1)(2x+3)) ]' / y^3 can you do it now ?

OpenStudy (anonymous):

Yep...thank you! :)

OpenStudy (anonymous):

yw. Im glad you can, this is a lot of work

terenzreignz (terenzreignz):

Even more work if logarithmic differentiation is not allowed :D